Show that ๐‘Š = {(๐‘ฅ, 4๐‘ฅ, 3๐‘ฅ) โˆˆ โ„2 |๐‘ฅ โˆˆ โ„} is a subspace of โ„ 3 . Also find a basis for subspace ๐‘ˆ of โ„ 3 which satisfies ๐‘Š โŠ• ๐‘ˆ = โ„3


1
Expert's answer
2022-05-10T11:49:32-0400

In order to prove that "W" is a subspace of "\mathbb{R}^3" we will use a subspace criteria. Namely, it is enough to prove that "u+w\in W" for any two vectors "u,w\in W" and "\alpha u\in W" for any "\alpha\in{\mathbb{R}}". Suppose that "u,w\in W". It means that "u=(x,4x,3x)" and "w=(y,4y,3y)" where "x,y\in{\mathbb{R}}." We obtain: "u+w=(x+y,4(x+y),3(x+y))". Denote "z:=x+y". We receive: "(z,4z,3z)\in W". Multiply a vector "u" by "\alpha": "\alpha u=(\alpha x,4\alpha x,3\alpha x)". Denote "\alpha x=a". We receive: "\alpha u=(a,4a,3a)\in W". Thus, the criteria holds. Remind that "W\oplus U=\{W+U|W\cap U=\empty\}". I.e., the direct sum consists of sums of vectors from "W" and "U". Consider a vector "(1,4,3)". it belongs to "W". The matrix "\left(\begin{array}{lll}1&4&3\\0&1&0\\0&0&1\end{array}\right)" is non-degenerate. The determinant is "1". It means that "(1,4,3)", "(0,1,0)","(0,0,1)" is a basis in "\mathbb{R}^3". Set "U=\{(0,x,y)|x,y\in{\mathbb{R}} \}". "W\cap U=\empty", since "(1,4,3),(0,1,0),(0,0,1)" are linearly independent. "W\oplus U=\mathbb{R}^3", since "(1,4,3),(0,1,0)(0,0,1)" is a basis.

Answer: "U=\{(0,x,y)|x,y\in{\mathbb{R}} \}". The basis of "U" is: "(0,1,0),(0,0,1)."